You are here: Foswiki>Commissioning Web>Com20120205 (10 Feb 2012, JuanGuerra)Edit Attach

UT 20120205 AO/PISCES technical night

Observer: DMiller (tucson), JCarlos (LBTO)
Telescope Operator: DHuerta (LBTO)
Telescope Support: JUrban(LBTO)
Instrument Support: DOfficer (LBTO)

Summary

-Test of the new preset modes to close the loop and apply offsets during the close loop. Several presets were sent with a good success in most all of them. The last preset, early in the morning failed during the autogain procedure, this was due to a bad seeing ~1.3" and the high background

-Check performances of the close loop.

-Check PSF quality in Pisces, the images showed a poor psf, this has to be investigated tomorrow (morning).

Details

First half was TCS Testing with TCS Build 2012B. TCS was reverted to BP14 for the second half.

Clouds passing. good seeing ~0.8" till the last hour where the seeing was about 1.3".

9:10 Opened chamber

Target in the ND holder plate.

We had a few minutes to check the ND holder plate using a diffuse source. The source is a piece of paper that cover the ND holder with a small hole in the center illuminated by a torch. A few images were taken and this test will be done tomorrow with more time to center and to fix the position for the ND holder plate.

Filter positions.

Images taken with 20 seconds exposure used to identify the filters posiitons. The table below shows the new positions..

filter band image
#9 Open pisces.20120205.0015.fits
0 dark pisces.20120205.0005,16.fits
1 J pisces.20120205.0007,17.fits
2 Ks pisces.20120205.0008,18.fits
3 H pisces.20120205.0009,19.fits
4 Br-gm pisces.20120205.0010,20.fits
5 FeII pisces.20120205.0011.fits
6 2.046um pisces.20120205.0012.fits
7 2.14um pisces.20120205.0013.fits
8 H2 pisces.20120205.0014.fits

long exposure images.

During the close dome we took a serie of images with light ON/OFF. The filter used was J. The test to check any possible vignetting in the optical path of Pisces.

image exptime(sec) light
pisces.20120205.0021,22 120 ON
pisces.20120205.0023,24 120 OFF
pisces.20120205.0025 240 ON
pisces.20120205.0026 240 OFF

Defocused images.

Preset to AO439.

Offset to center the star in the middle RA=10 DEC=5.

images taken:

offset RA,DEC (arcsec) image position
10 -2 pisces.20120205.0029.fits
offset 10,-3 image: pisces.20120205.0033.fits offset 10 -5

* Image of a defocus star, it is hard to identify a clear pupil image. defocused_20120205.png

10:06 Preset AO365

Offset along the detector. Images

10:45 First AdaptiveAce_Active preset

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=AO366
preset, OBJECT=AO366, MODE=ace, GS=1, RS=0, PA=37
auto_collimate
runao

  • 5.2 mag in H
  • Success. object in upper left quadrant, on the edge
  • bin1,
  • /newdata/pisces.20120205.00543fits J Saturated
  • offset, -10, -8, DETXY, object near the center
  • /newdata/pisces.20120205.0054.fits J Much Dimmer!!
  • offset, -2, 0, DETXY, object off edge now
  • /newdata/pisces.20120205.0055.fits J saturated

11:10 Second AdaptiveAce_Active preset

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=AO376
preset, OBJECT=AO376, MODE=ace, GS=4, RS=0, PA=37
auto_collimate
runao

  • 8.2 mag in H
  • bin 2, 990 Hz
  • Success. object not seen
  • /newdata/pisces.20120205.0056.fits J bject not seen
  • offset, -2, 0, DETXY
  • /newdata/pisces.20120205.0057.fits J object not seen
  • offset, -6, 0, DETXY
  • /newdata/pisces.20120205.0060.fits FeII object not seen
  • offset, -14, -8, DETXY to move to center
  • /newdata/pisces.20120205.0061.fits FeII object not seen
  • /newdata/pisces.20120205.0062.fits K object not seen
  • offset, -6, 0, DETXY
  • /newdata/pisces.20120205.0063.fits K (3 sec) 29000 peak counts
  • offset, -6, 2, DETXY
  • /newdata/pisces.20120205.0064.fits K (3 sec) not seen
  • offset, -6, 0, DETXY
  • /newdata/pisces.20120205.0063.fits K (3 sec) 29000 peak counts
  • offset, -6, -1, DETXY
  • /newdata/pisces.20120205.0065.fits K (3 sec) 5000 peak counts
  • offset, -18, 0, DETXY Lower left quadrant
  • /newdata/pisces.20120205.00568.fits K (3 sec) Saturated

12:02 third AdaptiveAce_Active preset R=14.2

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=AO476
preset, OBJECT=AO476, MODE=ace, GS=5, RS=0, PA=62
auto_collimate
runao

  • 9.6 mag in H
  • bin 2, 400 Hz
  • Success. object not seen
  • /newdata/pisces.20120205.0070.fits K Open loop, top left edge
  • offset, -6, 0, DETXY
  • /newdata/pisces.20120205.0073.fits K (3 sec) 1000 counts peak
  • offset, -14, -8, DETXY
  • /newdata/pisces.20120205.0074.fits K (3 sec) not seen
  • /newdata/pisces.20120205.0075.fits K (60 sec) not seen

12:29 fourth AdaptiveAce_Active preset R=15.7

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=AO457
preset, OBJECT=AO457, MODE=ace, GS=1, RS=0, PA=20
auto_collimate
runao

  • 10.9 in H
  • bin 3, 199 Hz
  • offset, -6, 0, DETXY
  • /newdata/pisces.20120205.0076.fits K (3 sec) not seen
  • /newdata/pisces.20120205.0077.fits K (60 sec) not seen
  • offset, -18, 0, DETXY
  • /newdata/pisces.20120205.0078.fits K (15 sec) 26000 peak counts
  • offset, -13, 0, DETXY
  • /newdata/pisces.20120205.0079.fits K (3 sec) 20 counts over background

12:29 fifth AdaptiveAce_Active preset R=15.9

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=AO384
preset, OBJECT=AO384, MODE=ace, GS=3, RS=0, PA=36
auto_collimate
runao

  • 10.6 H mag
  • bin 4, 311 Hz
  • Success.
  • /newdata/pisces.20120205.0080.fits K (15 sec) object not seen
  • offset, -18, 0, DETXY
  • /newdata/pisces.20120205.0081.fits K (15 sec) 7000 peak counts
  • offset, -18, -2, DETXY
  • /newdata/pisces.20120205.0082.fits K (15 sec) 2000 peak counts
  • /newdata/pisces.20120205.0083.fits K (15 sec) 2000 peak counts

13:08 fifth AdaptiveAce_Active preset R=13.7

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=M53
preset, OBJECT=M53, MODE=ace, GS=9, RS=0, PA=10
auto_collimate
runao

  • Dimm = 0.8" seeing
  • 11.11 H mag
  • bin 3, 399 Hz
  • /newdata/pisces.20120205.0084.fits K (15 sec) open loop
  • /newdata/pisces.20120205.0085.fits K (15 sec) 700 peak counts
  • offset, 0, -5, DETXY
  • /newdata/pisces.20120205.0086.fits K (10 sec) 2000 peak counts
  • /newdata/pisces.20120205.0086.fits K (10 sec) 2000 peak counts
  • /newdata/pisces.20120205.0088.fits K (60 sec) Open Loop

13:25 ADI preset

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=AO493
preset, OBJECT=AO493, MODE=acquire, GS=0, RS=0, PA=-10
pause
preset, OBJECT=AO493, MODE=ace, GS=6 RS=0, PA=-10, ROT=PARALLACTIC
auto_collimate
runao

  • rotator at 170. Need to calculate PA for PARALLACTIC mode rather than POSITION mode
  • Failed: Shell Ripped. Lost communication to slope computer. JC restarted slope computer process and BCU 39 restarted

13:35 Send just ADAPTIVEACE_TRACK

# Test of observing script
# 21Jun2011  DLM
header, PARTNER=ARIZONA, OBSERVER='Miller', OBJECT=AO493
preset, OBJECT=AO493, MODE=ace, GS=6 RS=0, PA=-20, ROT=PARALLACTIC
runao

  • Failed during Auto-gain procedure. Max skips reached, Stopping WFS, Stopping AdSec
  • ADI mode seemed to work

13:40 Closing

13:48 Pisces deactivated

AO loop performances.

With the problems due to the bas quality of the Pisces images we used teh elab tool to check the performances. Below is a table of the perfomarnce and a modal plot.

The quality of the loop close could be seen in the CCD47, where diffraction images could be seeing. Unfortunally the images were not saved.

After talking with the Arcetri Team, the note in the elab are not true, this is a wrong information from the elab. An issue track was activated: n#3939

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freqSorted ascending gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120205_123936 ??? 14.6 63 11 NaN 0.53 075552 3 66 200 2.7 1.2 1.0 6 41 ON -1.0 ? -1 -1 ONSKY NO  
20120205_124223 ??? 14.6 63 11 NaN 0.56 075552 3 66 200 2.7 1.2 1.0 6 40 ON -1.0 ? -1 -1 ONSKY NO  
20120205_124442 ??? 14.7 62 12 1.07 0.56 075552 3 66 200 2.7 1.2 1.0 6 39 ON -1.0 ? -1 -1 ONSKY NO  
20120205_124803 ??? 14.6 61 13 1.21 0.61 075552 3 66 200 2.7 1.2 1.0 6 40 ON -1.0 ? -1 -1 ONSKY NO  
20120205_125751 ??? 14.9 50 13 NaN 0.93 083227 4 36 312 3.4 1.5 0.8 6 40 ON -1.0 ? -1 -1 ONSKY NO  
20120205_125938 ??? 14.9 50 9 NaN 1.04 083227 4 36 312 3.4 1.5 0.8 6 39 ON -1.0 ? -1 -1 ONSKY NO  
20120205_130241 ??? 14.9 49 13 NaN 0.81 083227 4 36 312 3.4 1.5 0.8 6 40 ON -1.0 ? -1 -1 ONSKY NO  
20120205_130340 ??? 14.9 49 13 NaN 0.69 083227 4 36 312 3.4 1.5 0.8 6 41 ON -1.0 ? -1 -1 ONSKY NO  
20120205_130555 ??? 14.9 49 11 NaN 0.80 083227 4 36 312 3.4 1.5 0.8 6 40 ON -1.0 ? -1 -1 ONSKY NO  
20120205_130727 ??? 14.9 48 11 NaN 0.75 083227 4 36 312 3.4 1.5 0.8 6 39 ON -1.0 ? -1 -1 ONSKY NO  
20120205_121030 ??? 13.4 66 10 NaN 0.77 180534 2 145 400 2.4 1.8 1.0 3 27 ON -1.0 ? -1 -1 ONSKY NO  
20120205_121214 ??? 13.3 66 11 NaN 0.86 180534 2 145 400 2.4 1.8 1.0 3 30 ON -1.0 ? -1 -1 ONSKY NO  
20120205_121733 ??? 13.4 65 11 NaN 0.86 180534 2 145 400 2.4 1.8 1.0 3 28 ON -1.0 ? -1 -1 ONSKY NO  
20120205_122620 ??? 13.3 63 11 NaN 0.86 180534 2 145 400 2.4 1.8 1.0 3 30 ON -1.0 ? -1 -1 ONSKY NO  
20120205_131514 ??? 13.9 63 7 NaN 0.75 075552 3 66 400 2.7 0.9 0.9 7 40 ON -1.0 ? -1 -1 ONSKY NO - Pyramid modulation mismatch
20120205_131558 ??? 13.9 63 9 NaN 0.56 075552 3 66 400 2.7 0.9 0.9 7 40 ON -1.0 ? -1 -1 ONSKY NO - Pyramid modulation mismatch
20120205_131919 ??? 13.9 62 8 0.92 0.56 075552 3 66 400 2.7 0.9 0.9 7 40 ON -1.0 ? -1 -1 ONSKY NO - Pyramid modulation mismatch
20120205_132044 ??? 14.0 62 11 1.58 0.61 075552 3 66 400 2.7 0.9 0.9 7 37 ON -1.0 ? -1 -1 ONSKY NO - Pyramid modulation mismatch
20120205_105438 ??? 8.2 60 11 NaN 0.57 092604 1 400 990 0.9 0.7 0.7 3 313 ON -1.0 ? -1 -1 ONSKY NO - Pupils mismatch
20120205_105506 ??? 8.2 60 11 NaN 0.55 092604 1 400 990 0.9 0.7 0.7 3 314 ON -1.0 ? -1 -1 ONSKY NO - Pupils mismatch
20120205_110328 ??? 8.2 58 11 NaN 0.42 092604 1 400 990 0.9 0.7 0.7 3 319 ON -1.0 ? -1 -1 ONSKY NO - Pupils mismatch
20120205_112505 ??? 11.5 62 11 NaN 0.48 180534 2 145 991 1.4 0.5 0.4 3 64 ON -1.0 ? -1 -1 ONSKY NO  
20120205_114134 ??? 11.5 58 11 NaN 0.77 180534 2 145 991 1.4 0.5 0.4 3 65 ON -1.0 ? -1 -1 ONSKY NO  

Modal plot.

Condensation on the Pisces entrance window. Effect to the image quality

-- JuanGuerra - 05 Feb 2012
  • Modal plot bin 2:
    modalplot_b2.png

  • Image at close loop of the resulting PSF with condensation on the entrance window:
    consensation_20120205-0.png

  • Condensation, the image shows the creation of ice, this has been remove using the dry air hose:
    condensation_image_20120205.JPG

  • Defocuse bright star:
    defocused_20120205.png

  • Image offset to the center_with condensation:
    seeinglimited_offset_center.png
Topic attachments
I Attachment Action Size Date Who Comment
condensation_image_20120205.JPGJPG condensation_image_20120205.JPG manage 1 MB 10 Feb 2012 - 17:26 UnknownUser condensation
consensation_20120205-0.pngpng consensation_20120205-0.png manage 152 K 08 Feb 2012 - 12:45 UnknownUser Condensation on teh entrance window
defocused_20120205.pngpng defocused_20120205.png manage 1 MB 10 Feb 2012 - 17:28 UnknownUser Defocuse bright star
modalplot_b2.pngpng modalplot_b2.png manage 57 K 06 Feb 2012 - 06:13 UnknownUser Modal plot bin 2
seeinglimited_offset_center.pngpng seeinglimited_offset_center.png manage 406 K 10 Feb 2012 - 18:00 UnknownUser Image offset to the center_with condensation
Topic revision: r7 - 10 Feb 2012, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback